# JHU Applied Physics Laboratory, 11100 Johns Hopkins Rd, Laurel, MD 20723, USA

### v1

@@
The SPACE Program I: The featureless spectrum of HD 86226 c challenges sub-Neptune atmosphere trends
Sub-Neptune exoplanets are the most abundant type of planet known today. As they do not have a Solar System counterpart, many open questions exist about their composition and formation. Previous spectroscopic studies rule out aerosol-free hydrogen-helium-dominated atmospheres for many characterized sub-Neptunes but are inconclusive about their exact atmospheric compositions. Here we characterize the hot (Teq=1311 K) sub-Neptune HD 86226 c (R=2.2 , M=), which orbits its G-type host star on a 4 day orbit. The planet is located in a special part of the sub-Neptune parameter space: its high equilibrium temperature prohibits methane-based haze formation, increasing the chances for a clear atmosphere on this planet. We use HST data taken with WFC3 and STIS from the Sub-neptune Planetary Atmosphere Characterization Experiment (SPACE) Program to perform near-infrared (– ) transmission spectroscopy and ultraviolet characterization of the host star. We report a featureless transmission spectrum that is consistent within with a constant transit depth of ppm. The amplitude of this spectrum is only 0.01 scale heights for a H/He-dominated atmosphere, excluding a cloud-free solar-metallicity atmosphere on HD 86226 c with a confidence of . Based on an atmospheric retrieval analysis and forward models of cloud and haze formation, we find that the featureless spectrum could be due to a metal enrichment ( confidence lower limit) of a cloudless atmosphere, or alternatively silicate (MgSiO3), iron (Fe), or manganese sulfide (MnS) clouds. For these species, we perform a detailed investigation of cloud formation in high metallicity, high-temperature atmospheres. Our results highlight that HD 86226 c does not follow the aerosol trend of sub-Neptunes found by previous studies. Follow-up observations with the James Webb Space Telescope could determine whether this planet aligns with the recent detections of metal-enriched atmospheres or if it harbors a cloud species otherwise atypical for sub-Neptunes.
@@
Sub-Neptune exoplanets with a radius between and make up the most abundant planet class found by currently available surveys (e.g., Borucki et al., 2011; Batalha et al., 2013; Fulton et al., 2017). The absence of a Solar System analog highlights the importance of examining sub-Neptune exoplanets to understand this planet class. However, observing sub-Neptunes is challenging with current technology, as these small-radius planets only cover a fraction of their host stars’ light during the transit, leading to signal strengths of only a few hundred parts per million (ppm). Therefore, facilities with high sensitivities are needed to distinguish the planet signal from its star and various noise sources. The required precision is best met by space-based facilities such as the Hubble Space Telescope (HST) and the James Webb Space Telescope (JWST). But even these telescopes operate close to their limits to detect sub-Neptune planetary atmospheres.
@@
The few sub-Neptune atmospheres characterized today hint at very diverse atmospheric conditions on these planets. Their compositions range from hydrogen-helium dominated atmospheres on GJ 3470 b (Benneke et al., 2019), HD 3167 c (Guilluy et al., 2021), K2-18 b (Madhusudhan et al., 2023), and TOI-421 b (Davenport et al., 2025), to metal-rich, miscible atmospheres as detected on TOI-270 d (Benneke et al., 2024). The molecular inventory differs across these planets: CH4 is the dominant absorber in K2-18 b and TOI-270 d, H2O is seen in TOI-421 b, and SO2 absorption is detected in GJ 3470 b (Beatty et al., 2024) and is thought of as a consequence of disequilibrium photochemistry. Furthermore, many of those atmospheres possibly host aerosols, as detected in GJ 3470 b (Benneke et al., 2019; Beatty et al., 2024) and GJ 1214 b (Kreidberg et al., 2014a; Kempton et al., 2023).
@@
In contrast to the well-characterized planets, many of the observed sub-Neptune spectra are featureless or have muted spectral features. Examples of these include HD 97658 b (Knutson et al., 2014; Guo et al., 2020a), HD 106315 b (Kreidberg et al., 2022), TOI-836c (Wallack et al., 2024), GJ 1214 b (Kreidberg et al., 2014a; Kempton et al., 2023), and GJ 9827 d (Roy et al., 2023), where muted features are detected for the last two planets. These muted spectral features were predicted as a consequence of high mean molecular weight atmospheres (e.g., Miller-Ricci et al., 2009) or high-altitude aerosols (e.g., Benneke & Seager, 2013; Morley et al., 2013). Which of these effects dominates is still unclear for most of the observed planets. Trends in the spectra of Neptune-sized to sub-Neptune-sized planets suggest that there is a correlation between planet equilibrium temperature and atmospheric feature size (Crossfield & Kreidberg, 2017; Yu et al., 2021; Brande et al., 2024). Brande et al. (2024) attribute this trend to clouds, which rain out at low temperatures below . For temperatures above , a reduced formation of organic hazes may lead to clear atmospheres (e.g., Zahnle et al., 2009; Fortney et al., 2013; Crossfield & Kreidberg, 2017; Gao et al., 2020; Yu et al., 2021). The characterized sample of sub-Neptunes is small, and not all planets follow this trend. More systematic observations are needed to fully understand which environment favors which atmospheric conditions.
@@
The Sub-neptune Planetary Atmosphere Characterization Experiment (SPACE) with the HST is a multi-cycle treasury program that aims to conduct such a systematic survey (Kreidberg et al., 2022). It probes a sample of eight sub-Neptunes with near-infrared transmission spectroscopy using the Wide Field Camera 3 (WFC3111https://www.stsci.edu/hst/instrumentation/wfc3), and combines these data with ultraviolet (UV) stellar characterization using the Space Telescope Imaging Spectrograph (STIS222https://www.stsci.edu/hst/instrumentation/stis). As shown in the top panel of Fig. 3, the targets span a physically motivated grid across planet radius (2 – 3.5 ) and temperature (300 – 1400 ) to examine how the atmospheric conditions are altered with these parameters. Planet growth models (Zeng et al., 2019) predict the presence of atmospheres on all SPACE targets based on their radii and masses between and (see bottom panel of Fig. 3). The associated host stars range in spectral type between M and G, corresponding to the types with the most planet detections today. The experiment aims to reveal how sub-Neptune atmospheres are shaped by metal enrichment, disequilibrium chemistry, and aerosols. Furthermore, it will guide upcoming observations with JWST and identify sub-Neptunes with clear atmospheres.
@@
Here, we examine HD 86226 c, which is a sub-Neptune that orbits its G-type host star on a orbit. It has a high equilibrium temperature of (Teske et al., 2020), and its low mass of suggests the presence of an atmosphere (Teske et al., 2020) on the planet (see Fig. 3). These characteristics place HD 86226 c close to both the ‘radius gap’ (Fulton et al., 2017) and the ‘hot Neptune desert’ (Mazeh et al., 2016). Moreover, HD 86226 c has a long-period giant planet companion, HD 86226 b (Arriagada et al., 2010). These factors make HD 86226 c an interesting candidate for atmospheric characterization, as it could inform planet formation and evolution models. The G-type host star is both beneficial and hindering for the observations. While we are not expecting high stellar activity based on its estimated age of (Teske et al., 2020), the expected transit depth is only of order and, therefore, just on the lower edge of what we can observe with HST. However, in this sparsely populated part of the temperature-radius parameter space, HD 86226 c is one of the best available targets for this study.
@@
The hot temperatures on HD 86226 c suggest a high chance for a clear atmosphere. At these temperatures, CO will be the dominant carbon reservoir, instead of CH4. This hinders the formation of organic hazes (Zahnle et al., 2009; Fortney et al., 2013; Gao et al., 2020), which could mute the observed atmospheric features. Previous observations of HD 86226 c with the Transiting Exoplanet Survey Satellite (TESS) and the Magellan Telescopes (Kokori et al., 2023; Teske et al., 2020) constrained its orbital parameters and mass. Possible volatile-rich atmospheric compositions were recently predicted by Piette et al. (2023), but no spectroscopic observations of the planet’s atmosphere were conducted prior to this experiment.
@@
Section 2 provides details on the observations of HD 86226 c and its host star. The independent data reductions of the transit spectra and stellar data are presented in Sect. 3. The resulting light curves are analyzed in Sect. 4, where we also show the planetary spectrum. Forward models and retrievals of the planet spectrum are detailed and shown in Sect. 5. The implications of these results are discussed in Sect. 6 and Sect. 7 summarizes our findings.
@@ 2.1 HST observations with WFC3 and STIS
HD 86226 c was observed with HST as part of the Cycle 30 Large Treasury Program 17192 (PI: Laura Kreidberg). For the transit observations, we used the Wide Field Camera 3 (WFC3) with the G141 grism, spanning the wavelength between and . Individual exposures of the time series observations used the MULTIACCUM mode with three nondestructive readouts and a total integration time of . The spatial scanning mode was used to allow longer integration times before saturation by spreading the light across the detector (see e.g., Deming et al., 2013). The scan rate of 0.754 ”s-1 yields a scan height of 35.2” for the final readout, corresponding to approximately 290 pixels on the detector.
@@ 3.1 Transit reduction with PACMAN
For wavelength calibration, we used the stellar model grids of Castelli & Kurucz (2003). The model assumes the stellar parameters and , obtained from fits to the stellar spectral energy distribution with ARIADNE, which are described in Sect. 3.3. To obtain the calibration, we interpolated the stellar model and fitted it to the extracted observed spectrum. In this fit, we derived the shift in wavelength direction and allowed for a linear stretch in flux and wavelength.
@@
To correct this effect, we introduced a new systematics decorrelation parameter, the ‘row shift’, based on the position of the spectral trace. We independently assumed a linear correlation between flux and shift of spectral trace in spatial direction (row shift) for both scan directions. The row shift was determined for each exposure by summing the flux of the first non-destructive readout along the dispersion direction (columns) and comparing this spatial profile to the profile of the first exposure of all observations. This resulted in the row shift relative to the first exposure, which was used in the further analysis of the data.
@@ 3.2 Data reduction with Eureka!
Each stage in the pipeline is guided by ‘Eureka Control Files’ (ECF’s) and ‘Eureka Parameter Files’ (EPF’s), with the latter used to adjust transit model fit parameters. Eureka! currently provides template ECF’s for JWST’s MIRI, NIRCam, NIRSpec instruments, and HST’s WFC3 instrument. Further details can be found on the Eureka ReadTheDocs page.
@@
Eureka! also detected the non-negligible shift of the spectral trace in the scan direction on the detector (see Fig. 2). For seven of the eight transits analyzed with Eureka!, the shift in scan direction is greater than the recommended 0.2 pixel (peak-to-peak) threshold for precise time series observations (Stevenson & Fowler, 2019). Because each pixel has distinct sensitivity characteristics, spatial shifting of the spectrum introduces variability in the measured flux (Stevenson & Fowler, 2019). Correcting for these effects requires precise modeling of positional shifts that are dynamic and do not follow simple patterns. As such, modeling is currently not implemented in Eureka!, the positional dependence limits the precision that can be achieved with this pipeline. The limitations and capabilities of Eureka! for fitting the light curves of this data set will be discussed further in Sects. 4.2 and 4.3.
@@ 3.3 Stellar parameter estimation with ARIADNE
For the analysis of the planet data, it is essential to obtain precise parameters of the star, especially those needed to determine the best-fitting stellar models. We derived the stellar parameters based on fits to the spectral energy distribution (SED) with ARIADNE (Vines & Jenkins, 2022) code.
ARIADNE is an open-source Python package designed to automatically fit the spectral energy distribution (SED) of stars to different stellar atmosphere model grids. For HD 86226, we used the Phoenix v2 (Husser et al., 2013), BT-Settl (Hauschildt et al., 1999; Allard et al., 2012), Kurucz (1993), and Castelli & Kurucz (2003) model grids, which have been previously convolved with several broadband filters. ARIADNE operates under a Bayesian framework, where each grid is modeled using a Nested Sampling (Skilling, 2004, 2006; Higson et al., 2019) algorithm as implemented by dynesty (Speagle, 2020). The algorithm calculates the evidence of each model, while also computing the posterior distributions for the effective temperature, , [Fe/H], extinction in V, and radius of the star. After each grid was fitted, we combined the posterior distributions using Bayesian Model Averaging, which leverages the evidence of each model and accounts for model intrinsic uncertainties. We summarize the results from ARIANDE in Table 5.
@@ 3.4 STIS stellar spectral analysis
The STIS spectra were automatically calibrated with CalSTIS v.3.4.2 and retrieved from MAST. The automated extraction successfully identified the spectral trace in all three gratings, and no custom extractions were required. We detect multiple strong emission lines in the spectra and a clear continuum signal redwards of 1350. To splice the three spectra together, as well as the orders of the echelle E230M spectra, we retrieved the Hubble Advanced Spectral Products (HASP666https://archive.stsci.edu/missions-and-data/hst/hasp) complete coadd of the spectrum.
@@
The gap between the G140L and G230M spectra covering 1715–2274 Å and wavelengths 3118 Å were filled with a Phoenix model spectrum from the Lyon BT-Settl CIFIST 2011_2015 grid (Allard, 2016; Baraffe et al., 2015) using the stellar parameters found in Sect. 3.3. Uncertainties were estimated using models with the uncertainty. We find that the model is a good match to both sides of the gap in the UV spectrum, although the model diverges from the spectrum at wavelengths Å. This indicates that continuum contribution from the stellar chromosphere begins to dominate over the photosphere flux at those wavelengths.
Wavelengths below 1160 Å were filled using the FUV to Extreme ultraviolet (EUV) scaling relations from France et al. (2018). We used the scaling relationship for the Si iv 1400 Å doublet as it is detected with high S/N in the G140L spectrum. The line flux was measured by integrating the G140L spectrum over the region 1390–1410 Å then subtracting a flat continuum flux fitted to the surrounding spectrum. We find . Propagating both the error on the flux and on the scaling relationships, we find and . We generated a two-step spectrum for the region 90-1160 Å with these relationships, extending the higher wavelength band to meet the blue end of the STIS spectrum.
Our completed Spectral Energy Distribution (SED) is shown in Fig. 3 with the various components labeled. We compare the SED to that of the Sun: as might be expected from the stellar parameters, they are almost identical. The slight offsets between the spectra are within the uncertainties of the reconstructed spectrum.
Especially flux at wavelengths shorter than 1400 Å is strongly affected by stellar activity, in addition to the effective temperature. The close agreement of HD 86226 with the flux of the Sun therefore implies similar activity levels of both stars. This is also traced by the 25 day rotational period HD 86226 (Arriagada, 2011), similar to the solar rotation period at the equator.

### v2

@@
The SPACE Program. I. The featureless spectrum of HD 86226 c challenges sub-Neptune atmosphere trends
Sub-Neptune exoplanets are the most abundant type of planet known today. As they do not have a Solar System counterpart, many open questions exist about their composition and formation. Previous spectroscopic studies have ruled out aerosol-free hydrogen-helium dominated atmospheres for many characterized sub-Neptunes but are inconclusive about their exact atmospheric compositions. Here we characterize the hot (Teq=1311 K) sub-Neptune HD 86226 c (R=2.2 , M=), which orbits its G-type host star on a four-day orbit. The planet is located in a special part of the sub-Neptune parameter space: Its high equilibrium temperature prohibits methane-based haze formation, increasing the chances for a clear atmosphere on this planet. We used Hubble Space Telescope data taken with WFC3 and STIS from the Sub-neptune Planetary Atmosphere Characterization Experiment (SPACE) Program to perform near-infrared (– ) transmission spectroscopy and ultraviolet characterization of the host star. We report a featureless transmission spectrum that is consistent within with a constant transit depth of ppm. The amplitude of this spectrum is only 0.01 scale heights for a H/He-dominated atmosphere, excluding a cloud-free solar-metallicity atmosphere on HD 86226 c with a confidence of . Based on an atmospheric retrieval analysis and forward models of cloud and haze formation, we find that the featureless spectrum could be due to metal enrichment ( confidence lower limit) of a cloudless atmosphere, or silicate (MgSiO3), iron (Fe), or manganese sulfide (MnS) clouds. For these species, we performed a detailed investigation of cloud formation in high metallicity, high-temperature atmospheres. Our results highlight that HD 86226 c does not follow the aerosol trend of sub-Neptunes found by previous studies. Follow-up observations with the James Webb Space Telescope could determine whether this planet aligns with the recent detections of metal-enriched atmospheres or if it harbors a cloud species that is otherwise atypical for sub-Neptunes.
@@
Sub-Neptune exoplanets with a radius between and make up the most abundant planet class found by currently available surveys (e.g., Borucki et al., 2011; Batalha et al., 2013; Fulton et al., 2017). The absence of a Solar System analog highlights the importance of examining sub-Neptune exoplanets to understand this planet class. However, observing sub-Neptunes is challenging with current technology, as these small-radius planets only cover a fraction of their host stars’ light during their transit, leading to signal strengths of only a few hundred parts per million (ppm). Therefore, facilities with high sensitivities are needed to distinguish the planet signal from its star and various noise sources. The required precision is best met by space-based facilities such as the Hubble Space Telescope (HST) and the James Webb Space Telescope (JWST). But even these telescopes operate close to their limits to detect sub-Neptune planetary atmospheres.
@@
The few sub-Neptune atmospheres characterized today hint at very diverse atmospheric conditions on these planets. Their compositions range from hydrogen-helium dominated atmospheres on GJ 3470 b (Benneke et al., 2019), HD 3167 c (Guilluy et al., 2021), K2-18 b (Madhusudhan et al., 2023), and TOI-421 b (Davenport et al., 2025) to metal-rich, miscible atmospheres as detected on TOI-270 d (Benneke et al., 2024). The molecular inventory differs across these planets: CH4 is the dominant absorber in K2-18 b and TOI-270 d, H2O has been seen in TOI-421 b, and SO2 absorption has been detected in GJ 3470 b (Beatty et al., 2024) and is thought to be a consequence of disequilibrium photochemistry. Furthermore, many of those atmospheres possibly host aerosols, as detected in GJ 3470 b (Benneke et al., 2019; Beatty et al., 2024) and GJ 1214 b (Kreidberg et al., 2014a; Kempton et al., 2023).
@@
In contrast to the well-characterized planets, many of the observed sub-Neptune spectra are featureless or have muted spectral features. Examples of these include HD 97658 b (Knutson et al., 2014; Guo et al., 2020a), HD 106315 b (Kreidberg et al., 2022), TOI-836 c (Wallack et al., 2024), GJ 1214 b (Kreidberg et al., 2014a; Kempton et al., 2023; Schlawin et al., 2024), GJ 9827 d (Roy et al., 2023; Piaulet-Ghorayeb et al., 2024), and GJ 3090 b (Ahrer et al., 2025), where muted features are detected for the last three planets. These muted spectral features were predicted as a consequence of high mean molecular weight atmospheres (e.g., Miller-Ricci et al., 2009) or high-altitude aerosols (e.g., Benneke & Seager, 2013; Morley et al., 2013). Which of these effects dominates is still unclear for most of the observed planets. Trends in the spectra of Neptune-sized to sub-Neptune-sized planets suggest that there is a correlation between planet equilibrium temperature and atmospheric feature size (Crossfield & Kreidberg, 2017; Yu et al., 2021; Brande et al., 2024). Brande et al. (2024) attribute this trend to clouds, which rain out at low temperatures below . For temperatures above , a reduced formation of organic hazes may lead to clear atmospheres (e.g., Zahnle et al., 2009; Fortney et al., 2013; Crossfield & Kreidberg, 2017; Gao et al., 2020; Yu et al., 2021). The characterized sample of sub-Neptunes is small, and not all planets follow this trend. More systematic observations are needed to fully understand which environment favors which atmospheric conditions.
@@
The Sub-neptune Planetary Atmosphere Characterization Experiment (SPACE) with the HST is a multi-cycle treasury program that aims to conduct such a systematic survey (Kreidberg et al., 2022). It probes a sample of eight sub-Neptunes with near-infrared transmission spectroscopy using the Wide Field Camera 3 (WFC3111https://www.stsci.edu/hst/instrumentation/wfc3) and combines these data with ultraviolet (UV) stellar characterization using the Space Telescope Imaging Spectrograph (STIS222https://www.stsci.edu/hst/instrumentation/stis). As shown in the top panel of Fig. 3, the targets span a physically motivated grid across planet radius (2 – 3.5 ) and temperature (300 – 1400 ), which allows to examine how the atmospheric conditions are altered with these parameters. Planet growth models (Zeng et al., 2019) predict the presence of atmospheres on all SPACE targets based on their radii and masses between and (see bottom panel of Fig. 3). The associated host stars range in spectral type between M and G, corresponding to the types with the most planet detections today. The experiment aims to reveal how sub-Neptune atmospheres are shaped by metal enrichment, disequilibrium chemistry, and aerosols. Furthermore, it guides upcoming observations with JWST and identify sub-Neptunes with clear atmospheres.
@@
Here, we examine HD 86226 c, which is a sub-Neptune that orbits its G-type host star on a orbit. It has a high equilibrium temperature of (Teske et al., 2020), and its low mass of suggests the presence of an atmosphere (Teske et al., 2020) on the planet (see Fig. 3). These characteristics place HD 86226 c close to both the “radius gap” (Fulton et al., 2017) and the “hot Neptune desert” (Mazeh et al., 2016). Moreover, HD 86226 c has a long-period giant planet companion, HD 86226 b (Arriagada et al., 2010). These factors make HD 86226 c an interesting candidate for atmospheric characterization, as it could inform planet formation and evolution models. The G-type host star is both beneficial and hindering for the observations. While we do not expect high stellar activity based on its estimated age of (Teske et al., 2020), the expected transit depth is only on the order of and therefore just on the lower edge of what can be observed with HST. However, in this sparsely populated part of the temperature-radius parameter space, HD 86226 c is one of the best available targets for this study.
@@
The hot temperatures on HD 86226 c suggest a high chance for a clear atmosphere. At these temperatures, CO is the dominant carbon reservoir, instead of CH4. This hinders the formation of organic hazes (Zahnle et al., 2009; Fortney et al., 2013; Gao et al., 2020), which could mute the observed atmospheric features. Previous observations of HD 86226 c with the Transiting Exoplanet Survey Satellite (TESS) and the Magellan Telescopes (Kokori et al., 2023; Teske et al., 2020) constrained its orbital parameters and mass. Possible volatile-rich atmospheric compositions were recently predicted by Piette et al. (2023), but no spectroscopic observations of the planet’s atmosphere were conducted prior to this experiment.
@@
Section 2 provides details on the observations of HD 86226 c and its host star. The independent data reductions of the transit spectra and stellar data are presented in Sect. 3. The resulting light curves are analyzed in Sect. 4, where we also show the planetary spectrum. Forward models and retrievals of the planet spectrum are detailed and shown in Sect. 5. The implications of these results are discussed in Sect. 6, and we summarize our findings in Sect. 7.
@@ 2.1 HST observations with WFC3 and STIS
HD 86226 c was observed with HST as part of the Cycle 30 Large Treasury Program 17192 (PI: Laura Kreidberg). For the transit observations, we used WFC3 with the G141 grism, spanning the wavelength between and . Individual exposures of the time series observations used the MULTIACCUM mode with three nondestructive readouts and a total integration time of . The spatial scanning mode was used to allow for longer integration times before saturation by spreading the light across the detector (see e.g., Deming et al., 2013). The scan rate of 0.754 ”s-1 yields a scan height of 35.2” for the final readout, corresponding to approximately 290 pixels on the detector.
@@ 3.1 Transit reduction with PACMAN
For wavelength calibration, we used the stellar model grids of Castelli & Kurucz (2003). The model assumes the stellar parameters and , obtained from fits to the stellar spectral energy distribution (SED) with ARIADNE, which are described in Sect. 3.3. To obtain the calibration, we interpolated the stellar model and fitted it to the extracted observed spectrum. In this fit, we derived the shift in wavelength direction and allowed for a linear stretch in flux and wavelength.
@@
To correct this effect, we introduced a new systematics decorrelation parameter, the “row shift”, based on the position of the spectral trace. We independently assumed a linear correlation between flux and shift of spectral trace in spatial direction (row shift) for both scan directions. The row shift was determined for each exposure by summing the flux of the first non-destructive readout along the dispersion direction (columns) and comparing this spatial profile to the profile of the first exposure of all observations. This resulted in the row shift relative to the first exposure, which was used in the further analysis of the data.
@@ 3.2 Data reduction with Eureka!
Each stage in the pipeline is guided by “Eureka Control Files” (ECFs) and “Eureka Parameter Files” (EPFs), with the latter used to adjust transit model fit parameters. Eureka! currently provides template ECFs for JWST’s MIRI, NIRCam, NIRSpec instruments, and HST’s WFC3 instrument. Further details can be found on the Eureka ReadTheDocs page.
@@
Eureka! also detected the non-negligible shift of the spectral trace in the scan direction on the detector (see Fig. 2). For seven of the eight transits analyzed with Eureka!, the shift in scan direction is greater than the recommended 0.2 pixel (peak-to-peak) threshold for precise time series observations (Stevenson & Fowler, 2019). Because each pixel has distinct sensitivity characteristics, spatial shifting of the spectrum introduces variability in the measured flux (Stevenson & Fowler, 2019). Correcting for these effects requires precise modeling of positional shifts that are dynamic and do not follow simple patterns. As such, modeling is currently not implemented in Eureka!, the positional dependence limits the precision that can be achieved with this pipeline. The limitations and capabilities of Eureka! for fitting the light curves of this dataset is discussed further in Sects. 4.2 and 4.3.
@@ 3.3 Stellar parameter estimation with ARIADNE
For the analysis of the planet data, it is essential to obtain precise parameters of the star, especially those needed to determine the best-fitting stellar models. We derived the stellar parameters based on fits to the SED with ARIADNE (Vines & Jenkins, 2022) code.
ARIADNE is an open-source Python package designed to automatically fit the SED of stars to different stellar atmosphere model grids. For HD 86226, we used the Phoenix v2 (Husser et al., 2013), BT-Settl (Hauschildt et al., 1999; Allard et al., 2012), Kurucz (1993), and Castelli & Kurucz (2003) model grids, which were previously convolved with several broadband filters. ARIADNE operates under a Bayesian framework, where each grid is modeled using a Nested Sampling (Skilling, 2004, 2006; Higson et al., 2019) algorithm as implemented by dynesty (Speagle, 2020). The algorithm calculates the evidence of each model, while also computing the posterior distributions for the effective temperature, , [Fe/H], extinction in V, and radius of the star. After each grid was fitted, we combined the posterior distributions using Bayesian model averaging, which leverages the evidence of each model and accounts for model intrinsic uncertainties. We summarize the results from ARIADNE in Table 5.
@@ 3.4 STIS stellar spectral analysis
The STIS spectra were automatically calibrated with CalSTIS v.3.4.2 and retrieved from MAST. The automated extraction successfully identified the spectral trace in all three gratings, and no custom extractions were required. We detect multiple strong emission lines in the spectra and a clear continuum signal redward of 1350. To splice the three spectra together, as well as the orders of the echelle E230M spectra, we retrieved the Hubble Advanced Spectral Products (HASP666https://archive.stsci.edu/missions-and-data/hst/hasp) complete coadd of the spectrum.
@@
The gap between the G140L and G230M spectra covering 1715–2274 Å and wavelengths greater than 3118 Å were filled with a Phoenix model spectrum from the Lyon BT-Settl CIFIST 2011_2015 grid (Allard, 2016; Baraffe et al., 2015) using the stellar parameters from Sect. 3.3. Uncertainties were estimated using models with that account for the uncertainty in . We find that the model is a good match to both sides of the gap in the UV spectrum, although the model diverges from the spectrum at wavelengths smaller than Å. This indicates that the continuum contribution from the stellar chromosphere begins to dominate over the photosphere flux at those wavelengths.
Wavelengths below 1160 Å were filled using the FUV to extreme UV (EUV) scaling relations from France et al. (2018). We used the scaling relationship for the Si iv 1400 Å doublet as it is detected with high S/N in the G140L spectrum. The line flux was measured by integrating the G140L spectrum over the region 1390–1410 Å then subtracting a flat continuum flux fitted to the surrounding spectrum. We find . Propagating both the error on the flux and on the scaling relationships, we find and . We generated a two-step spectrum for the region 90-1160 Å with these relationships, extending the higher wavelength band to meet the blue end of the STIS spectrum.
Our completed SED is shown in Fig. 3 with the various components labeled. We compare the SED to that of the Sun. As might be expected from the stellar parameters, they are almost identical. The slight offsets between the spectra are within the uncertainties of the reconstructed spectrum.
The flux at wavelengths shorter than 1400 Å is strongly affected by stellar activity, in addition to the effective temperature. The close agreement of HD 86226 with the flux of the Sun therefore implies similar activity levels of both stars. This is also traced by the 25-day rotational period of HD 86226 (Arriagada, 2011), which is similar to the solar rotation period at the equator.